OpenStudy (anonymous):

Distance between 2 airports = 1,450 miles. Average speed of airplane = 350 mph. Travel time by bus after arrival at airport = 20 minutes. Total travel time =____ hours. 3.56 3.97 4.47

OpenStudy (kamibug):

Time equals distance divided by speed.

OpenStudy (anonymous):

\[Total~Time=\frac{ 1450 }{ 350 }+\frac{ 20 }{60 }=\frac{ 29 }{ 7 }+\frac{ 1 }{3 }=\frac{ 87+7 }{ 21 }=?\]